# MACTAS: Self-Attention-Based Module for Inter-Agent Communication in Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Communication is essential for the collective execution of complex tasks by human agents, motivating interest in communication mechanisms for multi-agent reinforcement learning (MARL). However, existing communication protocols in MARL are often complex and non-differentiable. In this work, we introduce a self-attention-based communication module that exchanges information between the agents in MARL. Our proposed approach is fully differentiable, allowing agents to learn to generate messages in a reward-driven manner. The module can be seamlessly integrated with any action-value function decomposition method and can be viewed as an extension of such decompositions. Notably, it includes a fixed number of trainable parameters, independent of the number of agents. Experimental results on the SMAC benchmark demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on several maps.
1 Introduction
@@
Multi-agent reinforcement learning (MARL) (Ning and Xie 2024; Du and Ding 2021; Nguyen, Nguyen, and Nahavandi 2020; Oroojlooy and Hajinezhad 2023; Wong et al. 2023) promises efficient control in large-scale systems through the collective learning of many agents in dynamic environments. Multiple agents interact within a shared environment to learn optimal policies. MARL is defined as the process in which agents optimize their behaviors based on the history of their interactions with the environment, i.e., on their experience. Research in this area aims to design methods that enable agents to learn quickly, i.e., from minimal experience. The challenges and complexities in using MARL for controlling multi-agent systems include coordination among agents, non-stationary environments due to evolving agent policies, and scalability issues. Advanced MARL approaches often involve sophisticated techniques such as decentralized training, communication protocols among agents, and hierarchical methods to manage multi-agent interactions as effectively as possible. Theoretical advancements and practical applications of MARL demonstrate its potential in diverse fields ranging from robotics (Orr and Dutta 2023) and autonomous driving (Yadav, Mishra, and Kim 2023; Schmidt et al. 2022; Althamary, Huang, and Lin 2019) to telecommunications (Li et al. 2022b; Feriani and Hossain 2021).
Figure 1: 3s5z_vs_3s6z scenario in the StarCraft Multi-Agent Challenge (SMAC).
@@
The most prominent research topics in MARL are the action-value function decompositions (Wang et al. 2021; Liu, Zhu, and Chen 2023), where agents share a joint reward according to their contribution, and inter-agent communication (Liu et al. 2023; Guo, Shi, and Fan 2023), where agents exchange information in order to coordinate their joint policy. The most challenging multi-agent environments are partially observable, usually via limiting agents’ information to their close proximity, and the successful achievement of their common goals is predicated upon coordination, which needs to involve some form of communication. In most research, the action-value function and inter-agent communication are issues addressed separately. Therefore, in this paper, we introduce communication into the very core of the action-value function decomposition.
@@
The prevailing paradigm in cooperative MARL is Centralized Training and Decentralized Execution (CTDE), where the training algorithm utilizes the global state, but agents can only use their own local observations. We argue that in most challenging MARL tasks, this is insufficient. Moreover, in most challenging MARL tasks, some form of inter-agent communication is a necessity, which can be provided relatively easily (e.g., through WiFi). Therefore, we introduce a communication protocol based on collecting information from each agent in a central entity and broadcasting the combined information to all agents. Because information aggregation is based on attention, our proposed architecture requires a constant number of weights plus an embedding for each agent. It is easy to apply using readily available and efficient implementations of the Transformer encoder layer.
@@
  • We verify MACTAS experimentally on several challenging maps from SMAC, giving excellent results and establishing state-of-the-art in some cases.
2 Related Work
@@
Single-Agent Reinforcement Learning (SARL) (Sutton and Barto 2018) addresses trial-and-error learning of sequential decision making under uncertainty. In Multi-Agent Reinforcement Learning (MARL) (Ning and Xie 2024), a group of agents learns to act in the same dynamic environment, collectively modifying its state. There are various settings for MARL, depending on whether the agents operate synchronously or asynchronously, whether they have common or adverse goals, and other circumstances. In this paper, we focus on the synchronous and cooperative MARL, with a single reward value for all agents at every time instance.
@@
MARL has been addressed by extending well-established SARL algorithms: Multi-Agent SAC (Wei et al. 2018), Multi-Agent PPO (Yu et al. 2022), Multi-Agent DDPG (Lowe et al. 2017).
@@
Generally, cooperative MARL adopts a centralized training with a decentralized execution (CTDE) paradigm, which suffers from the global action-value function, whose complexity grows exponentially with the number of agents (Kraemer and Banerjee 2016). Therefore, action-value function decomposition is a fundamental problem in MARL. Existing decomposed architectures include modules assigned to agents. These modules are fed with agents’ observations and their previous actions and produce agents’ contributions to the architecture’s output. The best action for the agent is the one that produces the highest contribution. In VDN (Sunehag et al. 2018), these contributions are just added. In QMIX (Rashid et al. 2018), a neural module, the mixing network, produces its weighted sum with positive weights. In Weighted QMIX (Rashid et al. 2020), the architecture is trained with a special weight assigned to the collective actions deemed the best. In QTRAN (Son et al. 2019), an additional module fed with all the actions of all agents corrects the weighted sum of the agents’ contributions. In Qatten (Yang et al. 2020), the mixing network is based on attention. In CDS (Li et al. 2021), a component of the mixing network optimizes the amount of exploration undertaken by the agents. In DVD (Li et al. 2022a), the internal state of the modules assigned to agents is also passed to the mixing networks. In SHAQ (Wang et al. 2022), the mixing network is enriched by several components based on Shapley value theory. In QPLEX (Wang et al. 2021), a carefully engineered duplex dueling of pairs of contributions of agents is introduced into the mixing network. In NA2Q (Liu, Zhu, and Chen 2023), a similar duplex dueling is performed with the use of attention.
@@
Another fundamental problem in the cooperative MARL is the coordination of agents’ activity with a certain communication protocol. A number of solutions have been proposed for that purpose, including CommNet (Sukhbaatar, Szlam, and Fergus 2016) which introduces differentiable communication by aggregating a learned broadcast vector from each agent, TarMAC (Das et al. 2019) uses signature-based soft attention to direct agent communication towards recipients, NDQ (Wang et al. 2019) applies two forms of regularization – mutual information maximization and message entropy minimization to optimize message content, GA-Comm (Liu et al. 2020) uses two-stage attention mechanism to limit number of messages exchanged between agents, IS (Kim, Park, and Sung 2021) encodes trajectory of future actions for each agent and shares them with others, in MAIC (Yuan et al. 2022) agents explicitly model their teammates and use it to bias their value function, DHCG (Liu et al. 2023) learns a dependency graph between agents and uses it to limit the number of communication edges, TEM (Guo, Shi, and Fan 2023) which uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, and ACE (Li et al. 2023) which proposes bidirectional action-dependent Q-learning.
@@
In addition, there are solutions in which the agents communicate in natural language, including: TWOSOME (Tan et al. 2024) which uses an LLM to evaluate the probabilities of joint action and applies LoRa for efficient PPO training, and Verco (Li et al. 2024) which separates the output of the LLM into communication and actions, allowing LLM agents to exchange interpretable verbal messages.
@@
In our setting, we consider communication methods that can be connected with an arbitrary action-value decomposition method. We also allow agents to send messages based only on their local information (current observation, historical observations, and historical actions). In this framework, the best results so far have been achieved by MAIC (Yuan et al. 2022).
3 Method
Figure 2: The proposed MACTAS architecture.
3.1 Problem definition
@@
In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (DecPOMDP) with communication. A DecPOMDP is defined by a tuple, $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{P},\mathcal{O},\mathcal{R}\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ is a team of agents, $\mathbb{S}$ is a space of environmental states, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is a space of team actions of the agents, $\mathcal{P}$ is the state transition probability, $\mathcal{O}$ is the observation function, and $\mathcal{R}$ is the reward function. At time $t=1,2,\dots$ each $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}^{i}(s_{t})$, of the environment state $s_{t}$. Agents may exchange messages with others through a certain channel that may generally refine these messages with some processing.
@@
Each agent maintains a representation, $\mathcal{I}^{i}_{t}$, of observations it has made, messages it has received, and actions it has taken. Based on thereof and a policy, $\pi^{i}$, the agent performs an action, $a^{i}_{t}\in\mathbb{A}^{i}$, $a^{i}_{t}\sim\pi^{i}(\cdot|\mathcal{I}^{i}_{t})$. The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t+1})$ is given collectively to all agents.
@@
The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ and the message exchange protocol for the team to expect in each environment state $s$ the highest sum of discounted rewards
$\mathcal{V}^{\pi}(s)=E\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big{|}s_{t}=s,\text{policy in use}=\pi\right),$
where $\gamma\in[0,1]$ is the discount factor.
@@
The goal of this paper is to establish a message exchange protocol that supports the above goal of learning.
3.2 MACTAS
@@
Our proposed Multi-Agent Communication via Transformer on Agents States (MACTAS) architecture is presented in Fig. 2. It is built on a decomposition of the action-value function. Each such decomposition includes the input modules assigned to the agents ($f^{i}$ assigned to the $i$-th agent for $i=1,\dots,n$ in Figure 2). These modules are fed with previous actions of these agents, $a^{i}_{t-1}$, and their current observations, $o^{i}_{t}$. They produce agents’ contributions, $\widehat{Q}^{i}_{t}$, to future discounted rewards and perhaps some other vectors (Li et al. 2022a). The outputs of the $f^{i}$ modules are fed to the mixing network that produces the value of the action-value function, $\widehat{Q}_{t}$. The input module $f^{i}$ usually comprises a feedforward part, denoted by MLP in Figure 2, a recurrent part, denoted by GRU, and the output feedforward part, again denoted by MLP.
@@
Intuitively, we understand the $f^{i}$ module as the consciousness of the $i$-th agent. Based on what the agent is aware of, it determines its contribution to future discounted rewards, depending on the action it currently takes. In order to give a global view on the current state to the agent’s consciousness, we introduce a communication mechanism. We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\mathbb{R}^{n_{h}}$ to the module $g$ where the states of consciousness of the agents are confronted. As a result, the $g$ module produces increments, $z^{i}_{t}\in\mathbb{R}^{n_{h}}$. The sum $z^{i}_{t}+h^{i}_{t}$ is then passed to the second MLP part of the $f^{i}$ module.
@@
  2. 2.
    It is a layered structure and the layers include self-attention (Vaswani et al. 2017).
Introducing $g$ satisfying the above requirements in their basic form is enough to outperform MARL without communication by a large margin. However, what follows from our experiments is that usually the best performing form of the $g$ module is a stack of Transformer (encoder) modules (Vaswani et al. 2017), with the following settings:
  1. 1.
@@
  3. 3.
    We use a separate Adam optimizer (Kingma and Ba 2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically it is RMSProp: Tieleman and Hinton 2012).
  4. 4.
    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module simply passes the input forward.
Good performance of the Transformer in this role is an expected result. This architecture has been specifically optimized to transform a set of embeddings into another set of the same size, which is exactly what we require in the $g$ module.
3.3 Implementation considerations
@@
Usually in multi-agent systems, reducing the communication capacity to $O(n)$ and offloading the communication processing burden to the IPU is very convenient. In these systems, the agents are naturally connected through a network and their computational capacity is limited (e.g., they are mobile and battery powered), contrary to that of the network infrastructure. Still, these are the agents that need to take on the processing of their observations and control.
In most inter-agent communication protocols, including MAIC (Yuan et al. 2022), all pairs of agents need to be able to communicate, which translates into $O(n^{2})$ messages. Technically, the parts of the agents that communicate with others could be moved to an IPU. That would require $O(n)$ messages to and from the IPU. Still, natural parallelization of MACTAS is its noticeable advantage.
@@
In the distributed implementation, it may be the case that some of the agents are out of reach of others and do not exchange information with them. That only means that self-attention in MACTAS is forced to assign zero weight to unavailable transmissions from those unheard agents.
3.4 Exploration
Learning in a multi-agent system with decentralized execution is hampered by agents’ policies falling into suboptimal Nash equilibria. Let us consider a typical MARL setting where agents use the $\epsilon$-greedy strategy with $\epsilon$ annealing and explore independently. On average, only $n\epsilon$ agents explore. In order to learn that a certain combination of agents’ actions is better than the current Nash equilibrium, at least two agents need to explore and try a better combination. However, this is unlikely. What is more likely is that only one agent explores, which leads to a combination worse than greedy. Note that this effect may be even stronger for a better approximation of the $Q$ function, as it leaves less room for accidental switches of the greedy action combinations.
To counteract this effect, we propose a combination of $\epsilon$-greedy and Boltzmann exploration. Namely, each agent with probability $\epsilon$ chooses independently a random action; however, with probability $1-\epsilon$ it does not act greedily, but chooses from the top $k$ of its actions (according to its local action-value function) using the Boltzmann strategy. Using $k=1$ is equivalent to the pure $\epsilon$-greedy policy. Selecting a small $k$ leads to exploring the region that is quite good according to the agent. We observe that $k=2$ works well in practice. The temperature of the Boltzmann exploration controls the amount of extra exploration that we want to add. We plan to extend this idea and other exploration strategies for MARL in future work.

### v2

@@ Abstract.
Communication is essential for the collective execution of complex tasks by human agents, motivating interest in communication mechanisms for multi-agent reinforcement learning (MARL). However, existing communication protocols in MARL are often complex and non-differentiable. In this work, we introduce a self-attention-based communication module that exchanges information between the agents in MARL. Our proposed approach is fully differentiable, allowing agents to learn to generate messages in a reward-driven manner. The module can be seamlessly integrated with any action-value function decomposition method and can be viewed as an extension of such decompositions. Notably, it includes a fixed number of trainable parameters, independent of the number of agents. Experimental results on the SMAC and SMACv2 benchmarks demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on a number of maps.
Key words and phrases:
Multi-Agent Reinforcement Learning, Cooperation in Multi-Agent Reinforcement Learning, Communication in Multi-Agent Reinforcement Learning
1. Introduction
@@
Multi-agent reinforcement learning (MARL) Ning and Xie (2024); Du and Ding (2021); Nguyen et al. (2020); Oroojlooy and Hajinezhad (2023); Wong et al. (2023) promises efficient control in large-scale systems through the collective learning of many agents in dynamic environments. Multiple agents interact within a shared environment to learn optimal policies. MARL is defined as the process in which agents optimize their behaviors based on the history of their interactions with the environment, i.e., on their experience. Research in this area aims to design methods that enable agents to learn quickly, i.e., from minimal experience. The challenges and complexities in using MARL for controlling multi-agent systems include coordination among agents, non-stationary environments due to evolving agent policies, and scalability issues. Advanced MARL approaches often involve sophisticated techniques, including decentralized training, communication protocols among agents, and hierarchical methods to manage multi-agent interactions as effectively as possible. Theoretical advancements and practical applications of MARL demonstrate its potential in diverse fields ranging from robotics Orr and Dutta (2023) and autonomous driving Yadav et al. (2023); Schmidt et al. (2022); Althamary et al. (2019) to telecommunications Li et al. (2022b); Feriani and Hossain (2021).
Figure 1. 3s5z_vs_3s6z scenario in the StarCraft Multi-Agent Challenge (SMAC).
@@
The most prominent research topics in MARL are the action-value function decompositions Wang et al. (2021); Liu et al. (2023b), where agents share a joint reward according to their contributions, and inter-agent communication Liu et al. (2023a); Guo et al. (2023), where agents exchange information in order to coordinate their joint policy. The most challenging multi-agent environments are partially observable, typically by limiting agents’ information to their close proximity, and the achievement of their common goals is predicated upon coordination, which needs to involve some form of communication. In most research, the action-value function and inter-agent communication are issues addressed separately. Therefore, in this paper, we introduce communication into the very core of the action-value function decomposition.
@@
The prevailing paradigm in cooperative MARL is Centralized Training and Decentralized Execution (CTDE), where the training algorithm utilizes the global state, but agents can only use their own local observations. We argue that in most challenging MARL tasks, this is insufficient. Moreover, in most challenging MARL tasks, some form of inter-agent communication is a necessity, which can be provided relatively easily (e.g., through WiFi). Therefore, we introduce a communication protocol that collects information from each agent in a central entity and broadcasts the combined information to all agents. Because information aggregation is based on attention, our proposed architecture requires a constant number of weights, plus an embedding, for each agent. It is easy to apply using readily available and efficient implementations of the Transformer encoder layer.
@@
  • We verify MACTAS experimentally on several challenging maps from SMACSamvelyan et al. (2019) and SMACv2 Ellis et al. (2023), giving excellent results and establishing state-of-the-art in some cases.
2. Related Work
@@
Single-Agent Reinforcement Learning (SARL) Sutton and Barto (2018) addresses trial-and-error learning in sequential decision-making under uncertainty. In Multi-Agent Reinforcement Learning (MARL) Ning and Xie (2024), a group of agents learns to act in the same dynamic environment, collectively modifying its state. There are various settings for MARL, depending on whether the agents operate synchronously or asynchronously, whether they have common or adverse goals, and other circumstances. In this paper, we focus on the synchronous and cooperative MARL, with a single reward value for all agents at every time instance.
@@
MARL has been addressed by extending well-established SARL algorithms, including Multi-Agent SAC Wei et al. (2018), Multi-Agent PPO Yu et al. (2022), and Multi-Agent DDPG Lowe et al. (2017).
@@
Methods of cooperative MARL often adopt a centralized training with decentralized execution (CTDE) paradigm, in which agents share information during training but act based solely on their own observations during execution. However, this paradigm suffers from the global action-value function, whose complexity grows exponentially with the number of agents Kraemer and Banerjee (2016). Therefore, action-value function decomposition is a fundamental problem in MARL. Existing decomposed architectures include modules assigned to agents. These modules are fed with agents’ observations and their previous actions and produce agents’ contributions to the architecture’s output. The best action for the agent is the one that makes the highest contribution. In VDN Sunehag et al. (2018), these contributions are just added. In QMIX Rashid et al. (2018), a neural module, the mixing network, produces its weighted sum with positive weights. In Weighted QMIX Rashid et al. (2020), the architecture is trained with a special weight assigned to the collective actions deemed the best. In QTRAN Son et al. (2019), an additional module fed with all the actions of all agents corrects the weighted sum of the agents’ contributions. In Qatten Yang et al. (2020), the mixing network is based on attention. In CDS Li et al. (2021), a component of the mixing network optimizes the amount of exploration undertaken by the agents. In DVD Li et al. (2022a), the internal state of the modules assigned to agents is also passed to the mixing networks. In SHAQ Wang et al. (2022), the mixing network is enriched by several components based on Shapley value theory. In QPLEX Wang et al. (2021), a carefully engineered duplex dueling of pairs of contributions of agents is introduced into the mixing network. In NA2Q Liu et al. (2023b), a similar duplex dueling is performed with the use of attention.
@@
Another fundamental problem in cooperative MARL is the coordination of agents’ activity for which a certain communication protocol is required. A number of solutions have been proposed for this purpose: CommNet Sukhbaatar et al. (2016) introduces differentiable communication by aggregating a learned broadcast vector from each agent, TarMAC Das et al. (2019) uses signature-based soft attention to direct agent communication towards recipients, NDQ Wang et al. (2019) applies two forms of regularization – mutual information maximization and message entropy minimization to optimize message content, GA-Comm Liu et al. (2020) uses two-stage attention mechanism to limit number of messages exchanged between agents, IS Kim et al. (2021) encodes trajectory of future actions for each agent and shares them with others, in MAIC Yuan et al. (2022) agents explicitly model their teammates and use it to bias their value function, DHCG Liu et al. (2023a) learns a dependency graph between agents and uses it to limit the number of communication edges, TEM Guo et al. (2023) uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, ACE Li et al. (2023) proposes bidirectional action-dependent Q-learning.
@@
In addition, there are solutions in which the agents communicate in natural language, including: TWOSOME Tan et al. (2024) which uses a large language model (LLM) to evaluate the probabilities of joint action and applies LoRa for efficient PPO training, and Verco Li et al. (2024) which separates the output of the LLM into communication and actions, allowing LLM agents to exchange interpretable verbal messages.
@@
We acknowledge various applications of attention or Transformer architectures in MARL algorithms, such as in Qatten mixing network, or in communication algorithms like NA2Q, GA-Comm, and TEM. None of the previous works has applied the Transformer architecture directly as a communication protocol. Our work fills this gap. We consider communication methods that can be coupled with an arbitrary action-value decomposition method. We also allow agents to send messages based only on their local information (current observation, historical observations, and historical actions). In this framework, the best results so far have been achieved by MAIC Yuan et al. (2022).
3. Method
Figure 2. The proposed MACTAS architecture.
3.1. Problem definition
@@
In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (Dec-POMDP) with communication. Dec-POMDP is defined by a tuple $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{R},\mathcal{P},\Omega,\mathcal{O},\gamma\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ represents the team of agents, $\mathbb{S}$ is the global environmental state space, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is the joint action space, $\mathcal{R}$ is the reward function, $\mathcal{P}$ is the state transition probability function, $\Omega=\Omega^{1}\times\dots\times\Omega^{n}$ is the joint observation space, $\mathcal{O}$ is the observation function and $\gamma$ is the discount factor. At time $t=1,2,\dots$ each $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}^{i}(s_{t}),o^{i}_{t}\in\Omega^{i}$, of the environment state $s_{t}\in\mathbb{S}$. Agents may exchange messages with others through a certain channel that may generally refine these messages with some processing.
@@
Each agent maintains a representation, $\mathcal{I}^{i}_{t}$, of observations it has made, messages it has received, and actions it has taken. Based on thereof and a policy, $\pi^{i}$, the agent performs an action, $a^{i}_{t}\in\mathbb{A}^{i}$, $a^{i}_{t}\sim\pi^{i}(\cdot|\mathcal{I}^{i}_{t})$. The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t},a_{t},s_{t+1})$ is given collectively to all agents.
@@
The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ to expect in each environment state $s$ the highest sum of discounted rewards
$\mathcal{V}^{\pi}(s)=\mathbb{E}\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big|s_{t}=s,\text{policy in use}=\pi\right),$
where $\gamma\in[0,1]$.
@@
The goal of this paper is to establish the message exchange protocol that supports the above goal of learning.
3.2. MACTAS
@@
Our proposed Multi-Agent Communication via Transformer on Agents States (MACTAS) architecture is presented in Fig. 2. It is built on a decomposition of the action-value function. Each such decomposition includes the input modules assigned to the agents ($f^{i}$ assigned to the $i$-th agent for $i=1,\dots,n$ in Figure 2). These modules are fed with previous actions of these agents, $a^{i}_{t-1}$, and their current observations, $o^{i}_{t}$. They produce agents’ contributions, $\widehat{Q}^{i}_{t}$, to future discounted rewards and perhaps some other vectors Li et al. (2022a). The outputs of the $f^{i}$ modules are fed to the mixing network that produces the value of the action-value function, $\widehat{Q}_{t}$. The input module $f^{i}$ usually comprises a feedforward part, denoted by MLP in Figure 2, a recurrent part, denoted by GRU, and the output feedforward part, again denoted by MLP.
@@
Intuitively, we understand the $f^{i}$ module as the consciousness of the $i$-th agent. Based on what the agent is aware of, it determines its contribution to future discounted rewards, depending on the action it currently takes. In order to give a global view on the current state to the agent’s consciousness, we introduce a communication mechanism. We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\real^{n_{h}}$ to the module $g$, where the states of consciousness of the agents are confronted. As a result, the $g$ module produces increments, $z^{i}_{t}\in\real^{n_{h}}$. The sum $z^{i}_{t}+h^{i}_{t}$ is then passed to the second MLP part of the $f^{i}$ module.
@@
  2. (2)
    It is a layered structure and the layers include self-attention Vaswani et al. (2017).
Introducing $g$ that satisfies the above requirements in their basic form is enough to outperform MARL without communication by a large margin. However, what follows from our experiments is that usually the best performing form of the $g$ module is a stack of Transformer (encoder) modules Vaswani et al. (2017), with the following settings:
  1. (1)
@@
  3. (3)
    We use a separate Adam optimizer Kingma and Ba (2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically it is RMSProp Tieleman and Hinton (2012)).
  4. (4)
    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module does not impact the state of consciousness of the agents.
A good performance of the Transformer in this role is an expected result. This architecture has been specifically optimized to transform a set of embeddings into another set of the same size, which is exactly what we require in the $g$ module.
3.3. Implementation considerations
@@
Usually in multi-agent systems, reducing the communication capacity to $O(n)$ and offloading the communication processing burden to the IPU is very convenient. In these systems, the agents are naturally connected through a network and their computational capacity is limited (e.g., they are mobile and battery-powered), contrary to that of the network infrastructure. Still, these are the agents that need to take on the processing of their observations and control.
In most inter-agent communication protocols, including MAIC Yuan et al. (2022), all pairs of agents need to be able to communicate, which translates into $O(n^{2})$ messages. Technically, the parts of the agents that communicate with others could be moved to an IPU. Communication through the IPU would require $O(n)$ messages from the agents to the IPU and back. Still, natural parallelization of MACTAS is its noticeable advantage.
@@
In the distributed implementation, it may happen that some of the agents are out of reach of others and do not exchange information with them. That only means that self-attention in MACTAS is forced to assign zero weight to unavailable transmissions from those unheard agents.
3.4. Exploration
Learning in a multi-agent system with decentralized execution is hampered by agents’ policies falling into suboptimal Nash equilibria. Let us consider a typical MARL setting where agents use the $\epsilon$-greedy strategy with $\epsilon$ annealing and explore independently. On average, only $n\epsilon$ agents explore. To learn that a certain combination of agents’ actions is better than the current Nash equilibrium, at least two agents need to explore and try a better combination. However, this happens rarely. What happens more often is that only one agent explores, which leads to a combination worse than greedy. Note that this effect may be even stronger for a better approximation of the $Q$ function, as it leaves less room for accidental switches of the greedy action combinations.
To counteract this effect, we propose a combination of $\epsilon$-greedy and Boltzmann exploration. Namely, each agent with probability $\epsilon$ chooses independently a random action; however, with probability $1-\epsilon$ it does not act greedily, but chooses from the top $k$ of its actions (according to its local action-value function) using the Boltzmann strategy. Using $k=1$ is equivalent to the pure $\epsilon$-greedy policy. Selecting a small $k$ leads to exploring a region that is quite good according to the agent. We observe that $k=2$ works well in practice. The temperature of the Boltzmann exploration controls the amount of extra exploration that we want to add.